Supplemental Operations Report – May 19th

Date: 5/19/2024
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Hab wall heater, solar observatory, perseverance parking brake & battery. EVA suit 9, Solar Power inverters slave 1&2, astronomy laptop dead

Power system:
Solar: Malfunctioning inverters still out for repairs.
SOC Last 24 hours: Max 100%; Min 90%; Avg 97.3%.
VDC Last 24 hours: Max 58.27V; Min 49.17V; Avg 53.86V.
Generator run time: 6139.7. Technician diagnosed problem as bad starter. They didn’t have a replacement so one was ordered. Ben installed on 5/16 and main generator is working again.

Propane Readings:
Station Tank: 77%
Director Tank: 71%
Intern Tank: 79%
Generator Tank: 74%

Hab Static Tank – 359 gallons
GreenHab – 0 gallons
Outpost tank – 370 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Sojourner rover used: No
Hours: 200.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: Perseverance is now charging only up to about 60%. Charging error appears from time to time. Rovers moved back to area in front of RAM now that the main generator is back online.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: No.
General notes and comments: Ignition problem persists in habcar. Can get it to start by resetting into park and trying to pull the shifter as high as possible.

Summary of Internet: All nominal.

EVA suits and radios:
Suits: Suit 9 has damaged visor and put aside until end of the season.
Comms: Nominal.

Campus wide inspection, if action taken, what and why: Solar Power still at 4.4kW. Main generator repaired and being used to charge rovers as needed.

Summary of Hab Operations: Wall heater is not functioning. Diagnostic is pending.
Summary of GreenHab Operations: Shut down for summer.
Summary of SciDome Operations: Inverters still out for repairs.

Summary of Observatories Operations: Lower shutter is decoupled and fastened by bungee cord. Astronomy laptop is dead.

Summary of RAM Operations: All nominal
Summary of Outpost Operations: All nominal
Summary of Health and Safety Issues: One crew member had what appears to have been a cold. They have recovered and all is nominal.

Copyright © The Mars Society. All rights reserved. | Main Site